# Correcting the Mythos of KL-Regularization: Direct Alignment without Overoptimization via $\chi^{2}$-Preference Optimization

### v2

@@ Appendix B Applying $\boldsymbol{\chi}$PO to the Token-Level MDP
Appendix C Sample Complexity Guarantees for $\boldsymbol{\chi^{2}}$-RLHF
Algorithm
Algorithm 3 $\chi^{2}$-RLHF
Theoretical guarantees
Assumption C.1.
Assumption C.2.
Below is our main sample complexity guarantee for $\chi^{2}$-RLHF. While it is stated for a fixed, $\beta$-dependent smoothing parameter for compactness, the most general version of this result (Theorem H.1) applies to general $\eta$.
Theorem C.1.
Let $\beta>0$ be given, and suppose Assumptions C.1 and C.2 hold any $\eta\in\left[0,\frac{\beta}{8R_{\mathsf{max}}}\right]$. With probability at least $1-\delta$, $\chi^{2}$-RLHF (Algorithm 3) produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
$J(\pi^{\star})-J(\widehat{\pi})\lesssim R_{\mathsf{max}}e^{2R_{\mathsf{max}}}\cdot\sqrt{\frac{\mathcal{C}^{\pi^{\star}}\log(|\mathcal{R}|/\delta)}{n}}+\beta\cdot\mathcal{C}^{\pi^{\star}}+\beta^{-1}\cdot\frac{R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\log(|\mathcal{R}|/\delta)}{n}+R_{\mathsf{max}}\sqrt{\frac{\log(|\Pi|/\delta)}{n_{\mathsf{x}}}}+\varepsilon_{\mathrm{opt}}.$
Above, we see that $\chi^{2}$-RLHF, like $\chi$PO, has sample complexity that scales only with the single-policy concentrability coefficient $\mathcal{C}^{\pi^{\star}}$, and holds for all comparator policies $\pi^{\star}$ simultaneously. Since the choice of $\beta$ induces a similar bias-overoptimization tradeoff in the first statement of Theorem C.1 as it did in Theorem 3.1 for $\chi$PO, we focus our discussion on the guarantee for a tuned choice of $\beta$ (Eq. 37). The first term in Eq. 37 accounts for the reward estimation error (2) and scales with $\mathcal{C}^{\pi^{\star}}$; as before, this accounts for how well rewards estimated from $\pi_{\mathsf{ref}}$ transfer to other candidate policies. The second term in Eq. 37 accounts for the statistical error from sampled contexts used in 4 for policy optimization. In particular, it is possible to drive this term to be much smaller than the first by using a larger unlabeled context dataset, which is typically far cheaper to acquire.
Computationally efficiency
Theorem C.1 bounds the sample complexity of $\chi^{2}$-RLHF under the assumption that we can solve 4 up to $\varepsilon_{\mathrm{opt}}$-accuracy. This is a purely computational problem, and in practice it can be solved using policy gradient methods such as PPO.
Comparison to $\boldsymbol{\chi}$PO
Unlike $\chi$PO (Theorem 3.1), Theorem C.1 has no dependence on the parameter $V_{\mathsf{max}}$ or quantities such as $\frac{\pi}{\pi_{\mathsf{ref}}}\leq\max_{\pi}\mathcal{C}_{\infty}^{\pi}$. We primarily attribute this to the fact that $\chi^{2}$-RLHF uses an explicit reward function class $\mathcal{R}$, and normalizing or clipping it to the reward range $R_{\mathsf{max}}$ is both natural and routinely done in practice (Shah et al., 2015; Christiano et al., 2017; Ouyang et al., 2022). In comparison, the implicit reward models induced by the policy class $\Pi$ in $\chi$PO can have larger range, and clipping the policy class in $\chi$PO directly, e.g., so that $|\beta\phi(\frac{\pi}{\pi_{\mathsf{ref}}})|$ is bounded, is misguided, because the policy class may lose realizability (Assumption 3.1). This is because $r^{\star}(x,a)=\beta\phi\Big{(}\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(a|x)}{\pi_{\mathsf{ref}}(a|x)}\Big{)}+Z_{\beta,r^{\star}}(x)$, and the normalization factor $Z_{\beta,r^{\star}}$ cannot be reasonably accounted for when clipping $\Pi$. While the $V_{\mathsf{max}}$ (Assumption 3.2) parameter involves pairs of action probabilities, and thereby sidesteps the normalization constant issue, it may not always be practical to modify $\Pi$ so that $V_{\mathsf{max}}$ is bounded, since this would require checking all pairs of each policy’s action probabilities.
However, using an explicit reward function class alone is not enough. As discussed previously, when we move from implicit to explicit $\chi^{2}$-regularization, incorporating the smoothing parameter $\eta$ in Eq. 35 is essential to avoid statistical errors due to policies with large density ratios when we approximate the $\chi^{2}$-regularizer empirically. A careful choice of $\eta=\beta/R_{\mathsf{max}}$ in Theorem C.1 balances the benefits of clipping against the bias it introduces. Without smoothing (i.e., $\eta=0$), a guarantee that depends on $\max_{\pi}\mathcal{C}_{\infty}^{\pi}$ for $\chi^{2}$-RLHF would be unavoidable, since the sample complexity must scale with the range of the problem, which grows with the magnitude of the regularizer. See Corollary H.2 in Appendix H for a guarantee in the case where $\eta=0$, which highlights this.
@@ Appendix D Preliminaries
Lemma D.1 (MLE for conditional density estimation (e.g., Wong and Shen (1995); de Geer (2000); Zhang (2006); Agarwal et al. (2020))).
@@
This section is organized as follows. First, in Section E.1, we analyze a more general version of $\chi$PO that mixes KL-regularization with $\chi^{2}$-regularization using a mixing parameter $\gamma\in(0,1]$, and present its sample complexity guarantee in Theorem E.1. $\chi$PO is a special case with $\gamma=1$, and Section E.2 shows (with a one-line proof) that Theorem 3.1 follows directly from Theorem E.1 with this parameter choice.
@@
Assumption E.1 (Policy realizability).
@@
Assumption E.2 (Bounded implicit rewards).
@@
Theorem E.1 (General version of Theorem 3.1).
Suppose Assumptions E.1 and E.2 hold for some $\beta>0$ and $\gamma\in(0,1]$. With probability at least $1-\delta$, the variant of $\chi$PO in Eq. 40 produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
@@
The bias-overoptimization tradeoffs induced by the choice of $\beta$ in Theorem E.1 are identical to those for Theorem 3.1 (and described there). Let us briefly discuss the influence of $\gamma$ on the sample complexity. We first observe that choice of $\gamma\in(0,1]$ changes the bound by only a small multiplicative factor, which implies that $\gamma$ can be arbitrarily small as long as it is positive. For the analysis, this is natural because the KL-divergence is dominated by the $\chi^{2}$-divergence, and, as discussed in Section 3.2, KL-regularization is only needed to enable the DPO-style reparameterization trick for Eq. 40 (in particular, the $\chi^{2}$-RLHF algorithm in Appendix C, which does not require reparameterization, obtains similar guarantees using pure $\chi^{2}$-regularization). It is worth noting, however, that the $\gamma$ parameter can implicitly influence the magnitude of $V_{\mathsf{max}}$, as well as the policy realizability condition. As such, practical consequences of this hyperparameter choice may not be fully captured by Theorem E.1.
Proof of Theorem E.1 Recall that the link function $\phi_{\gamma}$ induces a correspondence between policies in the class $\Pi$ and the implicit reward functions they induce (or, equivalently, between policies and the Bradley-Terry preference models they express). Our proof centers around the implicit reward model induced by the learned policy $\widehat{\pi}$,
@@
    (Lemma E.3) The reward model ${\widehat{r}}$ is an accurate estimate of $r^{\star}$ on the distribution of $\pi_{\mathsf{ref}}$. Moreover, we can transfer this guarantee to the distribution of any policy $\pi$ by paying a multiplicative $(1+2D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right))$-factor.
@@
    (Lemma E.2) $\widehat{\pi}$ maximizes the RLHF objective in Eq. 39 with reward model ${\widehat{r}}$, namely,
@@
The following lemma, which builds on a standard MLE generalization bound (Lemma D.1) bounds the error of ${\widehat{r}}$ under the action distribution induced by $\pi_{\mathsf{ref}}$. Recall that we use $\operatorname{\mathbb{E}}_{\pi,\pi^{\prime}}[\cdot]$ as shorthand for $\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x),b\sim\pi^{\prime}(\cdot\mid x)}[\cdot]$.
Lemma E.1.
Suppose Assumption E.1 holds. Then with probability at least $1-\delta$, the policy $\widehat{\pi}$ output by Eq. 40 satisfies
@@
Lemma E.1, along with all further supporting lemmas, is proven in the sequel. This result measures the error of ${\widehat{r}}$ using the clipped differences of rewards for pairs of actions $(x,a,b)$ drawn from $\pi_{\mathsf{ref}}$. Clipping the range of the implicit/explicit reward functions to $2R_{\mathsf{max}}$ ensures that the statistical error does not depend on $V_{\mathsf{max}}$. One minor but important detail in the proof is showing that Assumption E.1 implies $\mathcal{R}_{\Pi}$ includes the true reward function $r^{\star}$ up to an action-independent shift, so that the true preference model is realizable.
@@
This is a direct consequence of the result in Lemma E.2, which shows that an analogous property holds for general $f$-divergences. In particular, for any convex function $f$ and policy $\pi$, the policy $\pi$ is itself the optimal solution to the $f$-divergence-regularized RLHF objective under the implicit reward model induced by $\pi$ with the link function $f^{\prime}$.
Lemma E.2.
@@
Since $f_{\chi_{\textsf{mix}},\gamma}^{\prime}=\phi_{\gamma}=x+\gamma\log x$ for $\gamma>0$, clearly $0\not\in\mathrm{dom}(\phi_{\gamma})$. Further, under Assumption E.2, $\pi(a\mid{x})>0$ for all $\pi\in\Pi$ (otherwise $V_{\mathsf{max}}$ would be undefined), thus $\pi(a\mid{x})\in\mathrm{dom}(\phi_{\gamma})$ for all $(x,a)$. The claim in Eq. 44 then directly follows.
@@
To proceed, we will use condition on Lemma E.1 and use the event in this lemma to relate the estimated RLHF objective in Eq. 42 to the “true” RLHF objective that replaces ${\widehat{r}}$ with $r^{\star}$. An immediate challenge is that the RLHF objective in Eq. 42 must evaluate $\operatorname{\mathbb{E}}_{\pi}[{\widehat{r}}(x,a)]$ for all $\pi\in\Pi$, and accuracy under $\pi_{\mathsf{ref}}$ does not immediately imply that ${\widehat{r}}$ is accurate for other policies. The following bound quantifies the effects of this distribution shift using the $\chi^{2}$-divergence, and expresses how the estimation guarantee for ${\widehat{r}}$ in Lemma E.1 transfers to other policies $\pi$ of interest.
Lemma E.3.
Suppose Assumption 3.1 holds. Then for any $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$, under the event in Lemma E.1, we have
@@
where $\varepsilon_{\mathrm{stat}}^{2}$ is the off-policy estimation error defined in Lemma E.1.
It is worth noting that Lemma E.3 bounds the unclipped on-policy estimation error (on the LHS) in terms of the clipped off-policy estimation error, and in making this translation we pay for $V_{\mathsf{max}}$. As we will see shortly, working with the unclipped ${\widehat{r}}$ object is necessary for showing that Eq. 40 implicitly optimizes Eq. 42.
@@
In order to bound the reward estimation error terms in Eq. 45 using the guarantee we have previously established (Lemma E.3), we first center them using the return under the reference policy:
@@
where $\Delta^{\star}(x,a,b)\vcentcolon={}r^{\star}(x,a)-r^{\star}(x,b)$ and $\widehat{\Delta}(x,a,b)\vcentcolon={}{\widehat{r}}(x,a)-{\widehat{r}}(x,b)$. Substituting this identity back into the regret decomposition in Eq. 45, we apply Lemma E.3 with $\varepsilon_{\mathrm{stat}}^{2}\vcentcolon={}128R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\frac{\log(|\Pi|/\delta)}{n}$ (from Lemma E.1) to obtain
@@
Plugging in the expression for $\varepsilon_{\mathrm{stat}}^{2}$ results in the first statement of Theorem E.1.
@@
E.1.1 Proofs for Supporting Lemmas
Proof of Lemma E.1 Recall the reward-based MLE objective in Eq. 43,
@@ E.1.1 Proofs for Supporting Lemmas
which is equivalent to $r^{\star}$ up to an action-independent shift, namely, the normalization factor $\lambda^{\star}_{\beta,\gamma}$ in Lemma E.4. Since $\pi^{\star}_{\beta,\gamma}\in\Pi$ under Assumption E.1, we have $\widetilde{r}^{\star}\in\mathcal{R}_{\Pi}$, and for all $(x,a,b)\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}$, it holds that
@@
From this we conclude that $P_{\widetilde{r}^{\star}}=P_{r^{\star}}\in\mathcal{P}_{\Pi}$, and realizability is satisfied. Further, it is easy to see that $\mathcal{P}_{\Pi}$ contains only valid distributions. Thus, having satisfied the necessary preconditions, we can invoke Lemma D.1, which guarantees that with probability at least $1-\delta$, we have
@@
To conclude, we extract a bound on reward estimation error from this Hellinger distance bound by using Lemma E.5 with $R=V=2R_{\mathsf{max}}$, giving
@@
Proof of Lemma E.3 For a pair of policies $\pi,\pi^{\prime}$ and $p\geq 1$, we define the norm $\|\cdot\|_{p,\pi\times\pi^{\prime}}\vcentcolon={}\left(\operatorname{\mathbb{E}}_{\rho,a\sim\pi,b\sim\pi^{\prime}}[|\cdot|^{p}]\right)^{1/p}$. In addition, for notational compactness, we abbreviate $\widehat{\Delta}(x,a,b)\vcentcolon={}{\widehat{r}}(x,a)-{\widehat{r}}(x,b)$, and $\Delta^{\star}(x,a,b)\vcentcolon={}r^{\star}(x,a)-r^{\star}(x,b)$.
@@
This splits our bound into two terms. The first is the on-policy error of the clipped reward differences, and can be directly bounded by Lemma E.1 using a standard change-of-measure argument. The second expresses the error of translating the clipped estimates to the unclipped ones in our target bound. For the first term, using Cauchy-Schwarz gives
@@
In the second line we have used $\mathcal{C}^{\pi}=1+2D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)$ and the definition of $\varepsilon_{\mathrm{stat}}^{2}$ from Lemma E.1, and in the last line we use $V_{\mathsf{max}}\geq R_{\mathsf{max}}$.
Lemma E.4.
@@
Proof of Lemma E.4 It is easy to see that strong duality holds for Eq. 39, since it is convex and strictly feasible (e.g., for the policy $\pi_{\mathsf{ref}}$). Thus, the KKT conditions give the optimal primal and dual solutions.
@@
Next, for a function $g$ let $g^{-1}$ denote its left inverse, such that $g^{-1}(g(x))=x$. Because $\phi_{\gamma}$ is injective (see proof of Lemma E.2), it has a left inverse $(\phi_{\gamma})^{-1}$, and we can write
@@
Proof of Lemma E.5 We begin with the first statement, and write
@@
$D^{2}_{\mathsf{H}}\left(P_{z},P_{z^{\prime}}\right)=\sum_{y\in\{+1,-1\}}\frac{(P_{z}(y)-P_{z^{\prime}}(y))^{2}}{P_{z}(y)+P_{z^{\prime}}(y)}.$
@@
and therefore $(\sigma(z)-\sigma(z^{\prime}))^{2}\leq D^{2}_{\mathsf{H}}\left(P_{z},P_{z^{\prime}}\right)$. The result follows from taking the square root of both sides and combining with the first statement in the lemma. ∎
@@
Proof of Theorem 3.1 The policy optimization in 3 of Algorithm 1 is a special case of Eq. 40 with $\gamma=1$. As a result, Theorem 3.1 follows directly from Theorem E.1 when instantiated with $\gamma=1$. ∎
@@
Let $g_{x}(p)$ to denote the mixed divergence $\beta D_{f_{\chi_{\textsf{mix}}}}\left(p(x)\,\|\,\pi_{\mathsf{ref}}(x)\right)$. Then we have the following guarantee on regularized policy mirror descent (formal version of Lemma 6.1):
Lemma G.1.
@@
Then Assumption 6.2 indicates that $\overline{\pi}^{t+1}\in\Pi$ for all $t\in[T]$. In addition, by introducing Lagrangian multipliers into the above optimization problem and following similar arguments in the proof of Lemma E.4, we know
@@
Lemma G.2 (least squares regression).
@@
The proof of the above lemma is omitted. Applying Lemma G.2 to the least sqaures solution $\pi^{t+1}$, we have the following concentration lemma:
Lemma G.3 (concentration in optimization).
@@
In the following discussion, we use $\mathcal{E}_{1}$ to denote the event in Lemma G.3. Then under $\mathcal{E}_{1}$, by following the same arguments in the proof of Lemma E.3, we have the following bound on $\|f^{\beta,\eta}_{\overline{\pi}^{t+1},\pi^{t}}-f^{\beta,\eta}_{\pi^{t+1},\pi^{t}}\|_{1,\pi\times\pi_{\mathsf{ref}}}$:
@@
where $\rho\otimes\pi_{1}\otimes\pi_{2}$ denotes the joint distribution of $(x,a,b)$ where $x\sim\rho,a\sim\pi_{1}(x),b\sim\pi_{2}(x)$ for all $\pi_{1},\pi_{2}\in\Pi$. Applying the guarantee of least squares regression (Lemma G.2) to the least squares solution $\widehat{\ell}$, we have under Assumption 6.1, with probability at least $41-\delta/$ 4, the following event holds:
@@
G.3 Proofs for Supporting Lemmas
Proof of Lemma G.1 First for all $t\in[T],s\in\mathcal{S}$ and any policy $\pi\in\Pi_{C}$, we have

### v3

@@ Appendix E Preliminaries
Lemma E.1 (MLE for conditional density estimation (e.g., Wong and Shen (1995); de Geer (2000); Zhang (2006); Agarwal et al. (2020))).
@@
This section is organized as follows. First, in Section F.1, we analyze a more general version of $\chi$PO that mixes KL-regularization with $\chi^{2}$-regularization using a mixing parameter $\gamma\in(0,1]$, and present its sample complexity guarantee in Theorem F.1. $\chi$PO is a special case with $\gamma=1$, and Section F.2 shows (with a one-line proof) that Theorem 3.1 follows directly from Theorem F.1 with this parameter choice.
@@
Assumption F.1 (Policy realizability).
@@
Assumption F.2 (Bounded implicit rewards).
@@
Theorem F.1 (General version of Theorem 3.1).
Suppose Assumptions F.1 and F.2 hold for some $\beta>0$ and $\gamma\in(0,1]$. With probability at least $1-\delta$, the variant of $\chi$PO in Eq. 40 produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
@@
The bias-overoptimization tradeoffs induced by the choice of $\beta$ in Theorem F.1 are identical to those for Theorem 3.1 (and described there). Let us briefly discuss the influence of $\gamma$ on the sample complexity. We first observe that choice of $\gamma\in(0,1]$ changes the bound by only a small multiplicative factor, which implies that $\gamma$ can be arbitrarily small as long as it is positive. For the analysis, this is natural because the KL-divergence is dominated by the $\chi^{2}$-divergence, and, as discussed in Section 3.2, KL-regularization is only needed to enable the DPO-style reparameterization trick for Eq. 40 (in particular, the $\chi^{2}$-RLHF algorithm in Appendix B, which does not require reparameterization, obtains similar guarantees using pure $\chi^{2}$-regularization). It is worth noting, however, that the $\gamma$ parameter can implicitly influence the magnitude of $V_{\mathsf{max}}$, as well as the policy realizability condition. As such, practical consequences of this hyperparameter choice may not be fully captured by Theorem F.1.
Proof of Theorem F.1 Recall that the link function $\phi_{\gamma}$ induces a correspondence between policies in the class $\Pi$ and the implicit reward functions they induce (or, equivalently, between policies and the Bradley-Terry preference models they express). Our proof centers around the implicit reward model induced by the learned policy $\widehat{\pi}$,
@@
    (Lemma F.3) The reward model ${\widehat{r}}$ is an accurate estimate of $r^{\star}$ on the distribution of $\pi_{\mathsf{ref}}$. Moreover, we can transfer this guarantee to the distribution of any policy $\pi$ by paying a multiplicative $(1+2D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right))$-factor.
@@
    (Lemma F.2) $\widehat{\pi}$ maximizes the RLHF objective in Eq. 39 with reward model ${\widehat{r}}$, namely,
@@
The following lemma, which builds on a standard MLE generalization bound (Lemma E.1) bounds the error of ${\widehat{r}}$ under the action distribution induced by $\pi_{\mathsf{ref}}$. Recall that we use $\operatorname{\mathbb{E}}_{\pi,\pi^{\prime}}[\cdot]$ as shorthand for $\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x),b\sim\pi^{\prime}(\cdot\mid x)}[\cdot]$.
Lemma F.1.
Suppose Assumption F.1 holds. Then with probability at least $1-\delta$, the policy $\widehat{\pi}$ output by Eq. 40 satisfies
@@
Lemma F.1, along with all further supporting lemmas, is proven in the sequel. This result measures the error of ${\widehat{r}}$ using the clipped differences of rewards for pairs of actions $(x,a,b)$ drawn from $\pi_{\mathsf{ref}}$. Clipping the range of the implicit/explicit reward functions to $2R_{\mathsf{max}}$ ensures that the statistical error does not depend on $V_{\mathsf{max}}$. One minor but important detail in the proof is showing that Assumption F.1 implies $\mathcal{R}_{\Pi}$ includes the true reward function $r^{\star}$ up to an action-independent shift, so that the true preference model is realizable.
@@
This is a direct consequence of the result in Lemma F.2, which shows that an analogous property holds for general $f$-divergences. In particular, for any convex function $f$ and policy $\pi$, the policy $\pi$ is itself the optimal solution to the $f$-divergence-regularized RLHF objective under the implicit reward model induced by $\pi$ with the link function $f^{\prime}$.
Lemma F.2.
@@
Since $f_{\chi_{\textsf{mix}},\gamma}^{\prime}=\phi_{\gamma}=x+\gamma\log x$ for $\gamma>0$, clearly $0\not\in\mathrm{dom}(\phi_{\gamma})$. Further, under Assumption F.2, $\pi(a\mid{x})>0$ for all $\pi\in\Pi$ (otherwise $V_{\mathsf{max}}$ would be undefined), thus $\pi(a\mid{x})\in\mathrm{dom}(\phi_{\gamma})$ for all $(x,a)$. The claim in Eq. 44 then directly follows.
@@
To proceed, we will use condition on Lemma F.1 and use the event in this lemma to relate the estimated RLHF objective in Eq. 42 to the “true” RLHF objective that replaces ${\widehat{r}}$ with $r^{\star}$. An immediate challenge is that the RLHF objective in Eq. 42 must evaluate $\operatorname{\mathbb{E}}_{\pi}[{\widehat{r}}(x,a)]$ for all $\pi\in\Pi$, and accuracy under $\pi_{\mathsf{ref}}$ does not immediately imply that ${\widehat{r}}$ is accurate for other policies. The following bound quantifies the effects of this distribution shift using the $\chi^{2}$-divergence, and expresses how the estimation guarantee for ${\widehat{r}}$ in Lemma F.1 transfers to other policies $\pi$ of interest.
Lemma F.3.
Suppose Assumption 3.1 holds. Then for any $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$, under the event in Lemma F.1, we have
@@
where $\varepsilon_{\mathrm{stat}}^{2}$ is the off-policy estimation error defined in Lemma F.1.
It is worth noting that Lemma F.3 bounds the unclipped on-policy estimation error (on the LHS) in terms of the clipped off-policy estimation error, and in making this translation we pay for $V_{\mathsf{max}}$. As we will see shortly, working with the unclipped ${\widehat{r}}$ object is necessary for showing that Eq. 40 implicitly optimizes Eq. 42.
@@
In order to bound the reward estimation error terms in Eq. 45 using the guarantee we have previously established (Lemma F.3), we first center them using the return under the reference policy:
@@
where $\Delta^{\star}(x,a,b)\vcentcolon={}r^{\star}(x,a)-r^{\star}(x,b)$ and $\widehat{\Delta}(x,a,b)\vcentcolon={}{\widehat{r}}(x,a)-{\widehat{r}}(x,b)$. Substituting this identity back into the regret decomposition in Eq. 45, we apply Lemma F.3 with $\varepsilon_{\mathrm{stat}}^{2}\vcentcolon={}128R_{\mathsf{max}}^{2}e^{4R_{\mathsf{max}}}\frac{\log(|\Pi|/\delta)}{n}$ (from Lemma F.1) to obtain
@@
Plugging in the expression for $\varepsilon_{\mathrm{stat}}^{2}$ results in the first statement of Theorem F.1.
@@
F.1.1 Proofs for Supporting Lemmas
Proof of Lemma F.1 Recall the reward-based MLE objective in Eq. 43,
@@ F.1.1 Proofs for Supporting Lemmas
which is equivalent to $r^{\star}$ up to an action-independent shift, namely, the normalization factor $\lambda^{\star}_{\beta,\gamma}$ in Lemma F.4. Since $\pi^{\star}_{\beta,\gamma}\in\Pi$ under Assumption F.1, we have $\widetilde{r}^{\star}\in\mathcal{R}_{\Pi}$, and for all $(x,a,b)\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}$, it holds that
@@
From this we conclude that $P_{\widetilde{r}^{\star}}=P_{r^{\star}}\in\mathcal{P}_{\Pi}$, and realizability is satisfied. Further, it is easy to see that $\mathcal{P}_{\Pi}$ contains only valid distributions. Thus, having satisfied the necessary preconditions, we can invoke Lemma E.1, which guarantees that with probability at least $1-\delta$, we have
@@
To conclude, we extract a bound on reward estimation error from this Hellinger distance bound by using Lemma F.5 with $R=V=2R_{\mathsf{max}}$, giving
@@
Proof of Lemma F.3 For a pair of policies $\pi,\pi^{\prime}$ and $p\geq 1$, we define the norm $\|\cdot\|_{p,\pi\times\pi^{\prime}}\vcentcolon={}\left(\operatorname{\mathbb{E}}_{\rho,a\sim\pi,b\sim\pi^{\prime}}[|\cdot|^{p}]\right)^{1/p}$. In addition, for notational compactness, we abbreviate $\widehat{\Delta}(x,a,b)\vcentcolon={}{\widehat{r}}(x,a)-{\widehat{r}}(x,b)$, and $\Delta^{\star}(x,a,b)\vcentcolon={}r^{\star}(x,a)-r^{\star}(x,b)$.
@@
This splits our bound into two terms. The first is the on-policy error of the clipped reward differences, and can be directly bounded by Lemma F.1 using a standard change-of-measure argument. The second expresses the error of translating the clipped estimates to the unclipped ones in our target bound. For the first term, using Cauchy-Schwarz gives
@@
In the second line we have used $\mathcal{C}^{\pi}=1+2D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)$ and the definition of $\varepsilon_{\mathrm{stat}}^{2}$ from Lemma F.1, and in the last line we use $V_{\mathsf{max}}\geq R_{\mathsf{max}}$.
Lemma F.4.
@@
Proof of Lemma F.4 It is easy to see that strong duality holds for Eq. 39, since it is convex and strictly feasible (e.g., for the policy $\pi_{\mathsf{ref}}$). Thus, the KKT conditions give the optimal primal and dual solutions.
@@
Next, for a function $g$ let $g^{-1}$ denote its left inverse, such that $g^{-1}(g(x))=x$. Because $\phi_{\gamma}$ is injective (see proof of Lemma F.2), it has a left inverse $(\phi_{\gamma})^{-1}$, and we can write
@@
Proof of Lemma F.5 We begin with the first statement, and write
@@
$2D^{2}_{\mathsf{H}}\left(P_{z},P_{z^{\prime}}\right)\geq\sum_{y\in\{+1,-1\}}\frac{(P_{z}(y)-P_{z^{\prime}}(y))^{2}}{P_{z}(y)+P_{z^{\prime}}(y)}.$
@@
and therefore $(\sigma(z)-\sigma(z^{\prime}))^{2}\leq 2D^{2}_{\mathsf{H}}\left(P_{z},P_{z^{\prime}}\right)$. The result follows from taking the square root of both sides and combining with the first statement in the lemma. ∎
@@
Proof of Theorem 3.1 The policy optimization in 3 of Algorithm 1 is a special case of Eq. 40 with $\gamma=1$. As a result, Theorem 3.1 follows directly from Theorem F.1 when instantiated with $\gamma=1$. ∎
@@
Let $g_{x}(p)$ to denote the mixed divergence $\beta D_{f_{\chi_{\textsf{mix}}}}\left(p(x)\,\|\,\pi_{\mathsf{ref}}(x)\right)$. Then we have the following guarantee on regularized policy mirror descent (formal version of LABEL:lem:md-general-informal):
Lemma H.1.
@@
Then Assumption 7.2 indicates that $\overline{\pi}^{t+1}\in\Pi$ for all $t\in[T]$. In addition, by introducing Lagrangian multipliers into the above optimization problem and following similar arguments in the proof of Lemma F.4, we know
@@
Lemma H.2 (least squares regression).
@@
The proof of the above lemma is omitted. Applying Lemma H.2 to the least sqaures solution $\pi^{t+1}$, we have the following concentration lemma:
Lemma H.3 (concentration in optimization).
@@
In the following discussion, we use $\mathcal{E}_{1}$ to denote the event in Lemma H.3. Then under $\mathcal{E}_{1}$, by following the same arguments in the proof of Lemma F.3, we have the following bound on $\|f^{\beta,\eta}_{\overline{\pi}^{t+1},\pi^{t}}-f^{\beta,\eta}_{\pi^{t+1},\pi^{t}}\|_{1,\pi\times\pi_{\mathsf{ref}}}$:
@@
where $\rho\otimes\pi_{1}\otimes\pi_{2}$ denotes the joint distribution of $(x,a,b)$ where $x\sim\rho,a\sim\pi_{1}(x),b\sim\pi_{2}(x)$ for all $\pi_{1},\pi_{2}\in\Pi$. Applying the guarantee of least squares regression (Lemma H.2) to the least squares solution $\widehat{\ell}$, we have under Assumption 7.1, with probability at least $41-\delta/$ 4, the following event holds:
@@
H.3 Proofs for Supporting Lemmas
Proof of Lemma H.1 First for all $t\in[T],s\in\mathcal{S}$ and any policy $\pi\in\Pi_{C}$, we have